Circular No. 8396

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ea, 2004eb, AND 2004ec Three apparent supernovae have been reported from their appearance

on unfiltered CCD images: SNe 2004ea and 2004eb by S. Park and W. Li (LOSS; cf. IAUC 8395), and SN 2004ec by T. Boles (cf. IAUC 8392).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004ea Aug. 29.51 4^{h}16^{m}13.^{s}20 −16^{o}45′20′′.5 18.2 9′′.2 E, 20′′.3 S
2004eb Aug. 29.15 17 28 24.30 +57 32 45.7 17.1 3′′.4 E, 0′′.9 N
2004ec Aug. 29.949 17 08 04.96 +26 22 44.4 17.5 2′′.2 W, 0′′.8 N

Additional magnitudes from the respective discoverers: SN 2004ea in MCG –03-11-19, Feb. 13.26 UT, [20.0; Aug. 30.45, 18.3. SN 2004eb in the eastern component of the interacting galaxy pair NGC 6387, Aug. 25.17, [19.0; 30.18, 17.0. SN 2004ec in UGC 10717, May 27, [19.5; June 27, [19.5; Aug. 30.855, 17.3. Boles adds that nothing is visible at the location of SN 2004ec on Digitized Sky Survey images form 1990 (blue) and 1996 (red).

RS OPHIUCHI B. E. Schaefer, Louisiana State University, writes on the discovery of

a nova eruption of RS Oph in early 1907, as evidenced by the presence of a post-eruption dip in the light curve on archival photographic plates at the Harvard College Observatory. From 1907 May 3 to Sept. 4, the magnitude was at 12.6 < B < 13.0 on eight plates, while RS Oph was at 11.2 < B < 12.2 before 1906 Oct. 20 (the seasonal gap due to proximity to the sun is from 1906 Oct. 21 to 1907 May 2) and after 1907 Sept. 18 (with three plates from Sept. 18–28, just before the start of the seasonal gap). Such dips (with durations of 0.4–0.8 year) are seen only after all eruptions of this recurrent nova, so RS Oph must have had an eruption in early 1907 when it was hidden by the sun.

V1647 ORIONIS C. Aspin, Gemini Observatory; and B. Reipurth, University of Hawaii,

report that an acquisition-camera image, obtained in morning twilight on Aug. 29 UT through an R-band filter at the Gemini 8-m telescope on Mauna Kea, shows that V1647 Ori (cf. IAUC 8284, 8354) — now emerging from conjunction with the sun — continues to be in its elevated state, and that McNeil’s nebula remains bright. The duration of the outburst so far exceeds nine months.

2004 August 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green